OpenStudy (anonymous):

Kali left school and traveled toward her friend's house at an average speed of 40 km/h. Matt left one hour later and traveled in the opposite direction with an average speed of 50 km/h. Find the number of hours Matt needs to travel before they are 400 km apart.

OpenStudy (anonymous):

At the time of Matt's departure, Kali is already 40km away. From then on the rate of change will be ( 40km + 50km ) per hour. There the equation would be 40 + 90n, where n is the number of hours passed. If 40 + 90n = 400 Rearrange the equation to find n. 400-40 = 90n n = 360/90 n = 4 hours